OpenStudy (anonymous):

If the zeros of f(x) are x = -1 and x = 2, then the zeros of f(x/2) are x =

OpenStudy (anonymous):

hey do u hve d solution for dis question coz i m nt sure abt mine?

OpenStudy (anonymous):

umm i got -1/2 and 1

OpenStudy (anonymous):

i wasn't sure if the equation called for dividing the x values or the y values

OpenStudy (anonymous):

i hve a diff interpretation bt m nt sure. after expressin f(x) as a quadratic expression usin d roots,i plugged in x/2.

OpenStudy (anonymous):

that kind of confuses me. so you found an equation using the roots and then how did you plug in x/2? what answer did you get? (sorry i dont have a calculator with me right now)

OpenStudy (anonymous):

f(x)=(x+1)(x-2) f(x/2)=(x/2+1)(x/2-2) x/2+1=0 x=-2 nd x/2-2=0 x=4

OpenStudy (anonymous):

ohhh i understand, that makes sense now. Thanks so much

OpenStudy (anonymous):

glad 2 b ov sum help!!